Let the production for company A, Q= 20K^0.5L^0.5
a)The firm is currently producing 200 units of output per period using input rate of L= 4 and K= 25. Is this efficient input combination?Why or why not? If not determine the efficient input combination for producing an output rate of 200.
1
Expert's answer
2020-06-08T11:06:33-0400
"200=20 \times 4^{0.5}\times 25^{0.5}"


"200=20 \times 2 \times5"

"200=200"

This combination is rational.